find the x-intercept and the y-intercept for the graph of the equation.
x+2y=10

Answer by stanbon(75887)

find the x-intercept and the y-intercept for the graph of the equation.
x+2y=10
----
x-intercept = ?
Let y = 0 and solve for "x":
x + 2*0 = 10
x = 10
---
x-intercept: (10,0)
=============================
y-intercept = ?
Let x = 0 and solve for "y":
0 + 2y = 10
y = 5
---
y-intercept: (0,5)
=======================
Cheers,
Stan H.